OpenStudy (anonymous):

John began his job making $25 the first day. After that he was paid $6.75 per hour.Make a T-table for the equation for this, using the x-values: 0, 20, 40. Type in the ordered pairs you get as your answer, in this form: (0, number), (20, number), (40, number).

OpenStudy (anonymous):

y=6.75x+25 then plug in the x values

OpenStudy (anonymous):

so then i would just have y= 25+6.25*0 y=25+6.25*20 y=25+6.25*40

OpenStudy (anonymous):

yesh

OpenStudy (anonymous):

okay but what is a T table

OpenStudy (anonymous):

then add like terms. oh a T table! lol here i will draw one for ju

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):



OpenStudy (anonymous):

x values go on the left, y values go on the right!

OpenStudy (anonymous):

what is the x and y values? the 0 20 40 r the x right ??

OpenStudy (anonymous):